Suppose a wholesaler buys new disposable phones from your factory for $30 per phone, and wants to make a 30% margin on the selling price. How much must the wholesaler sell each phone?
Answer #1

Let the selling price be P

Profit margin is 30% of the selling price which means the margin will be 0.3P

Margin is given by P - MC

0.3P = P - 30

P = 30/0.7 = 42.86

Therefore the selling price of the phone should be 42.86.